# 70% Size, 100% Accuracy: Lossless LLM Compression for Efficient GPU Inference via Dynamic-Length Float

### v1

@@ 3.3 Dynamic-Length Float: A Lossless LLM Compression Framework for Efficient GPU Inference
40:                  $\mathbf{Outputs}[\mathbf{ThreadOutputPos}[t]]\leftarrow$
[…]
42:                      (Exponentbitwise_left_shift $(\mathrm{Exponent}\;\texttt{bitwise\_left\_shift}\;7)\;\texttt{bitwise\_or}\;\mathrm{Mantissa}$
43:                  $\mathrm{BitOffset}\leftarrow\mathrm{BitOffset}+\mathbf{CodeLengths}[\mathrm{Exponent}]$
44:                  $\mathbf{ThreadOutputPos}[t]\leftarrow\mathbf{ThreadOutputPos}[t]+1$
45:              end while
46:         end for
47:     end for
48:end procedure
3.3.1 Efficient Decoding with Compact LUTs
Huffman codes can be efficiently decoded using a lookup table (LUT)–based approach [35]. We construct a lookup table, denoted as $\mathbf{LUT}$, of size $2^{L}$, where $L$ is the maximum bit length of any Huffman code in the codebook. Each entry in $\mathbf{LUT}$ maps an $L$-bit sequence to the decoded symbol whose Huffman code is a prefix of that sequence. Due to the prefix-free property of Huffman codes, this mapping is guaranteed to be unambiguous.
To decode, we read the next $L$ bits from the encoded bitstream and use them as an index into $\mathbf{LUT}$ to retrieve the next decoded symbol. To determine the exact number of bits consumed in the decoding process (i.e., the length of the matched Huffman code), we use a second lookup table, $\mathbf{CodeLengths}$, which maps each symbol to the length of its Huffman code. We then advance the bitstream by that length and repeat the process to decode subsequent symbols.
For decoding the exponents of DFloat11, we constrain the maximum code length $L$ to 32 bits for each model, allowing us to read 32 bits from the encoded bitstream in each decoding step. Empirically, we observe that most models naturally produce maximum code lengths $L\leq 32$, or only slightly above this threshold. For models where $L>32$, we enforce the length constraint by reducing the frequencies of the least common exponents to 1 and rebuilding the Huffman tree. This results in a more balanced structure in the tail of the Huffman tree, assigning equal-length codes to the rarest exponents and reducing the maximum code length to 32 bits.
However, when $L=32$, a direct lookup table would require $2^{32}\approx 4.29$ billion entries, which is prohibitively memory-intensive. To address this, we propose partitioning the monolithic $\mathbf{LUT}$ into four disjoint and memory-efficient lookup tables. Specifically, we replace the $2^{32}$-entry array with four $2^{8}$-entry arrays: $\mathbf{LUT}_{1}$, $\mathbf{LUT}_{2}$, $\mathbf{LUT}_{3}$, and $\mathbf{LUT}_{4}$. Each entry in these tables consumes a single byte, resulting in a total memory footprint of $10244\times 2^{8}=$ 1024 bytes. The $\mathbf{CodeLengths}$ table adds another $2562^{8}=$ 256 bytes, bringing the total to 1280 bytes. This fits in GPU SRAM and enables fast access.
Decoding a 32-bit encoded sequence using the four compact LUTs proceeds as follows. We start by reading the first byte (8 bits) of the sequence and using it as an index into $\mathbf{LUT}_{1}$. If the corresponding Huffman code has a length of 8 bits or fewer, $\mathbf{LUT}_{1}$ returns a valid decoded symbol, and the decoding step is complete. Otherwise, it returns a reserved value $R$, indicating that the code has not yet been fully resolved. In that case, we read the next byte from the sequence and use it to index into $\mathbf{LUT}_{2}$. This process continues with $\mathbf{LUT}_{3}$ and $\mathbf{LUT}_{4}$ as needed, until a non-$R$ value is returned, signaling that the Huffman code has been successfully decoded. The first non-$R$ value returned is the decoded symbol corresponding to a prefix of the 32-bit encoded sequence.
While this approach significantly reduces memory usage, it introduces potential ambiguity. Specifically, if two different Huffman codes with distinct prefixes both map to the reserved value $R$ within the same LUT, the remaining bits may no longer be sufficient to unambiguously identify the correct decoded symbol in the subsequent LUTs. Fortunately, such conflicts are rare in practice and can be resolved. Due to the highly imbalanced distribution of exponent frequencies, the resulting Huffman trees are often skewed. As a result, each compact LUT typically maps at most one index to the value $R$, thereby avoiding ambiguity. In the rare cases where ambiguity does arise, it can be resolved by slightly adjusting the frequency distribution. Specifically, we increase the frequency of the more common exponent associated with the conflicting Huffman code and rebuild the Huffman tree. This shortens the code for that exponent, eliminating the conflict and ensuring that each compact LUT maps at most one index to the reserved value $R$.
3.3.2 Two-Phase Kernel and Lightweight Auxiliary Variables
To enable massively parallel decoding of entropy-coded exponents in DFloat11, we assign each thread a fixed number of bytes from the encoded sequence to process. However, this approach introduces two key challenges: 1. Because Huffman codes have variable bit widths and are tightly packed, the starting bit position for each thread to begin decoding is unclear. 2. Except for the first thread, the index of the elements being decoded is unknown, making it difficult to determine the correct output location for storing the results.
To address the first issue, we use a gap array [35] to determine the starting bit position for each thread. The gap array $\mathbf{Gaps}$ contains one entry per thread, and each entry specifies the bit offset of the first valid Huffman code relative to the thread’s assigned starting byte. Since the maximum code length is 32 bits, each offset lies in the range $[0,31]$. To maintain memory efficiency, we encode each entry using 5 bits.
To address the second issue, the most straightforward approach is to maintain an array that stores the output position of the first decoded element for each thread. However, this incurs substantial storage overhead. Given the large size of weight matrices, each output position must be represented as a 32-bit integer. Since decoding typically involves tens of thousands of threads per weight matrix, storing these positions results in significant memory consumption. This overhead undermines the compression ratio achieved by DFloat11.
To reduce storage overhead, we store the output position only for the first element of each thread block, rather than for every individual thread. Since each block typically contains hundreds to thousands of threads, this reduces the overhead from one 32-bit integer per thread to one per block, resulting in a negligible memory overhead.
To make decoding possible with block-level output positions, we adopt a two-phase kernel design. In the first phase, all threads within a block decode their assigned portions of the encoded sequence in parallel, but without writing any output to global memory. Instead, each thread counts how many elements it will decode. After this pass, we synchronize the threads within each block and compute the output position for every thread by calculating prefix sums over the counts, starting from the known output position of the block. In the second phase, each thread re-decodes the same portion of the encoded sequence, this time writing the decoded results to the HBM at the correct output positions. To avoid redundant memory read access to HBM during the two passes, we load the encoded exponents into SRAM. The pseudocode for the two-phase kernel is presented in Algorithm 1.
3.3.3 Transformer-Block-Level Decompression
We now have a complete method for decompressing entropy-coded exponents in a massively parallel manner. The weights of the LLM are stored in the DFloat11 format, accompanied by lightweight auxiliary data: thread-level gap offsets and block-level output positions, which determine the read and write locations for each thread. Both the compressed weights and auxiliary variables reside entirely on the GPU during inference. When a weight matrix is needed for matrix multiplication, it is decompressed on-the-fly into the original BFloat16 format. Once the matrix multiplication is complete, the BFloat16 matrix is immediately discarded to conserve GPU memory.
In practice, decompressing a single weight matrix often underutilizes GPU resources due to its relatively small size. In the DFloat11 decompression kernel, we set the number of bytes processed per thread to $n=8$, the number of threads per block to $T=256$, and the number of thread blocks to $B=\lceil\frac{|\mathbf{EncodedExponent}|}{nT}\rceil$, where $|\mathbf{EncodedExponent}|$ is the total number of bytes in the encoded exponents. As the size of the DFloat11 weights increases, more thread blocks are utilized, resulting in higher decompression throughput. This effect is illustrated in Figure 6, which shows that decompression throughput improves significantly with larger matrix sizes. To capitalize on this, we propose batching the decompression of multiple matrices together to improve throughput and hide latency. Specifically, we batch the decompression of all DFloat11 weight matrices within a single transformer block. Before executing any computation in a transformer block, we first decompress all its associated weights.
This technique significantly reduces decompression latency and improves overall inference efficiency. The latency breakdown for DFloat11-compressed Llama-3.1-8B-Instruct [11] across different batch sizes is shown in Figure 5.
4 Experiments
Table 2: Lossless compression statistics for various models. Model sizes are shown before and after compression.
Model Original $\to$ Losslessly Compressed Compression Ratio Compressed Bit Width
Llama-3.1-8B-Instruct 16.06 GB $\to$ 11.24 GB 69.98% 11.20
Meta-Llama-3-8B 16.06 GB $\to$ 11.26 GB 70.12% 11.22
Llama-3.3-70B-Instruct 141.11 GB $\to$ 96.14 GB 68.13% 10.90
Llama-3.1-405B-Instruct 811.71 GB $\to$ 551.22 GB 67.91% 10.87
Qwen2.5-14B-Instruct 29.54 GB $\to$ 20.81 GB 70.44% 11.27
Qwen2.5-32B-Instruct 65.53 GB $\to$ 45.53 GB 69.48% 11.12
QwQ-32B 65.53 GB $\to$ 45.53 GB 69.48% 11.12
Mistral-Nemo-Instruct-2407 24.50 GB $\to$ 17.04 GB 69.55% 11.13
Mistral-Small-24B-Instruct-2501 47.14 GB $\to$ 32.30 GB 68.52% 10.96
Codestral-22B-v0.1 44.49 GB $\to$ 30.24 GB 67.96% 10.87
gemma-2-9b-it 20.32 GB $\to$ 14.59 GB 71.81% 11.49
gemma-3-12b-it 25.55 GB $\to$ 18.21 GB 71.27% 11.40
gemma-3-27b-it 56.84 GB $\to$ 39.95 GB 70.28% 11.25
DeepSeek-R1-Distill-Qwen-7B 15.23 GB $\to$ 10.73 GB 70.48% 11.28
DeepSeek-R1-Distill-Llama-8B 16.06 GB $\to$ 11.23 GB 69.95% 11.19
In this section, we empirically evaluate the effectiveness of DF11 compression and its inference efficiency on GPUs. We compress a wide selection of recent LLMs from their original BFloat16 representation into our proposed DF11 format and report the resulting compression factors. We then compare the inference performance of the DF11-compressed models running on different GPUs with that of uncompressed models. Finally, we conduct an ablation study to analyze the effects of compression.
Software and Hardware We implement the DF11 decompression kernel in CUDA and C++, and integrate it into the Transformers [33] inference framework. We evaluate the inference efficiency of our losslessly compressed models against their uncompressed counterparts. For the uncompressed baseline, we use the HuggingFace Accelerate framework to support CPU offloading and multi-GPU inference. To assess the performance of the DF11 kernel across different hardware configurations, we run experiments on multiple machines with varying GPU and CPU setups. The hardware specifications for all experimental machines are provided in Table 4 in the Appendix.
Table 3: Comparison of accuracy and perplexity for the original (BF16) and losslessly compressed (DF11) models on different benchmarks. DF11 compression results in absolutely no loss in accuracy or perplexity.
Accuracy Perplexity
Model Data Type MMLU TruthfulQA WikiText C4
Llama-3.1-8B-Instruct BF16 68.010 $\pm$ 0.375 36.965 $\pm$ 1.690 8.649 21.677
DF11 (Ours) 68.010 $\pm$ 0.375 36.965 $\pm$ 1.690 8.649 21.677
Qwen2.5-14B-Instruct BF16 78.885 $\pm$ 0.330 51.775 $\pm$ 1.749 6.661 23.093
DF11 (Ours) 78.885 $\pm$ 0.330 51.775 $\pm$ 1.749 6.661 23.093
4.1 Results
Figure 3: Comparison of average latency and throughput for token decoding between the original (BF16) models and their losslessly compressed (DF11) counterparts. Portions of the BF16 models are offloaded to the CPU due to GPU memory constraints.
Figure 4: Comparison of GPU memory consumption between the original (BF16) models and their losslessly compressed (DF11) counterparts. The DF11 models support 5.33–13.17$\times$ longer context lengths by allowing more GPU memory to be used for storing the KV cache. “O.O.M.” means out of memory.
4.1.1 DF11 Compresses LLMs to 70% Size
Table 2 presents the compression factors of DF11 for a wide selection of recent LLMs. Specifically, we apply compression to all linear projection layers by converting their parameters from BF16 to DF11. The models we compress include LLaMA 3/3.1/3.3 [11], Qwen 2.5 [36], QwQ [31], Mistral Nemo/Small/Codestral [29, 30, 28], Gemma 2/3 [27, 26], and DeepSeek R1 Distilled [12]. DF11 achieves approximately 70% compression across all models, corresponding to an effective bit width of around 11 bits.
4.1.2 Accuracy and Perplexity Evaluations Confirm DF11 Is Perfectly Lossless
We verify the lossless property of DF11 compression through a series of accuracy and perplexity evaluations on standard benchmarks. Evaluations are conducted using lm_evaluation_harness [9], reporting accuracy on MMLU [15] and TruthfulQA [22], and word-level perplexity on WikiText [24] and C4 [25]. The results are shown in Table 3. As demonstrated, the compressed models achieve identical accuracy and perplexity to their original BF16 counterparts. To further validate the losslessness property, we compare the decompressed BF16 weight matrices from DF11 against the originals for each model in Table 2, confirming exact bit-level equivalence.
4.1.3 DF11 Outperforms CPU Offloading in Inference Efficiency
We compare the inference efficiency of DF11 and BF16 models across different hardware platforms. The uncompressed BF16 models exceed the memory limits of a single GPU, whereas the losslessly compressed DF11 models fit within those limits. For the BF16 models, we retain most of the model and the computation in the GPU while offloading some components and their associated computations to the CPU. To measure latency and throughput, we first perform a warm-up run by processing 100 tokens. In the actual evaluation, we decode 100 tokens starting from an empty prompt, using varying batch sizes. Each configuration is run five times, and we report the average latency and throughput across these runs. The results across different models, GPUs, and batch sizes are presented in Figure 3. As shown, DF11 models consistently outperform the BF16 models with CPU offloading, achieving 1.85–38.83$\times$ lower latency or higher throughput. For comparison using multiple GPUs, Figure 9 in the Appendix shows the performance of DF11 models running on a single GPU versus BF16 models running on two GPUs.
4.1.4 Memory Savings from DF11 Enable Longer Generation Lengths
The memory savings provided by DF11 compression not only reduce the number of GPUs required for inference but also enable longer generation lengths. During inference, the KV cache grows lienarly with the number of decoded tokens and quickly becomes a bottleneck for GPU memory. In Figure 4, we show the GPU memory consumption of DF11 and BF16 models during inference with a batch size of 1, as the number of decoded tokens increases. As shown, DF11 compression significantly extends the token generation length, allowing 5.33–13.17$\times$ more tokens to be decoded before reaching the GPU memory limit compared to BF16 models.
Figure 5: Comparison of latency breakdown for DFloat11 and BFloat16 Llama-3.1-8B-Instruct during GPU inference for different token batch sizes, using one A100-40GB GPU.
Figure 6: Throughput (top two) and latency (bottom two) comparisons between transferring BFloat16 matrices from CPU to GPU and decompressing the same matrices using the NVIDIA nvCOMP ANS library and our proposed DFloat11 kernel, across matrix sizes and GPU types.
4.2 Ablation Study
4.2.1 Latency Breakdown Shows Decompression Overhead Is Amortized at Larger Batch Sizes
We compare the latency breakdown of Llama-3.1-8B-Instruct in BF16 and DF11 formats using varying token batch sizes on a single A100-40GB GPU. For each configuration, we measure the latency of each component during the forward pass over 10 runs and report the average in Figure 5. Compared to the original model, the DF11-compressed version introduces additional latency due to the decompression of transformer blocks and the language modeling head. This decompression adds a constant overhead that is independent of the token batch size. As a result, increasing the batch size amortizes the decompression cost, leading to a substantially smaller gap in overall inference time.
4.2.2 DF11 Decompression Is Significantly Faster Than CPU-to-GPU Transfer and ANS
We compare the latency and throughput of the DF11 decompression kernel against two baselines: CPU-to-GPU transfer and ANS (Asymmetric Numeral System) decompression [5] from the NVIDIA nvCOMP library [1], using weight matrices of varying sizes. The motivation is to evaluate two strategies for memory-constrained inference: offloading weight matrices to CPU memory and transferring them to the GPU as needed, or storing compressed matrices on the GPU and decompressing them on demand. For this study, we use the weight matrix of the language modeling head from Llama-3.1-8B-Instruct and slice it into varying sizes. The results are shown in Figure 6. As illustrated, DF11 decompression is significantly more efficient than both baselines, achieving up to 24.87$\times$ higher throughput or lower latency than CPU-to-GPU transfer, and up to 15.12$\times$ faster than NVIDIA nvCOMP decompression. In addition to speed, DF11 also provides better compression ratio than nvCOMP, reducing model size by approximately 70% compared to around 78% for nvCOMP. Notably, DF11 decompression throughput improves with larger matrix sizes due to increased GPU thread utilization.
5 Related Works
Data Formats for Model Weights
Compressed data formats typically aim for lower bit-widths. For example, FP8—which comes in both E4M3 (4 exponent bits, 3 mantissa bits, plus 1 sign bit) and E5M2 configurations—has seen reasonable adoption in LLM training and development. Integer formats like INT8 have also been well explored, as in LLM.int8() [2] and its following works. Formats with a stronger emphasis on efficiency, such as FP4, INT4, NF4 [3], and AF4 [39], use only 4 bits. In this work, we primarily focus on formats with $\geq$8 bits, as benchmark literature [37, 10, 23] often suggests that 8-bit quantization results in negligible performance drop—though we show in Section 2 that this claim is likely skewed due to evaluation selectiveness and benchmark limitations.
Lossless Model Compression
While lossy model compression techniques such as pruning and quantization [6, 21, 7] have received widespread attention, lossless model compression remains a relatively underexplored area. Upon careful investigation, we identified roughly four prior works that have made meaningful efforts in this space. Deep Compression [13] is a foundational work, applying Huffman coding [17] to quantized CNN models and achieving an additional $\sim$22compression gain for model checkpoints. ZipNN [16] extended this idea to language models, comparing its results to classic lossless compression tools such as zlib [4] and zstd [^3] https://github.com/facebook/zstd and demonstrated superior compression gains. However, this line of work is limited in that its efficiency gains only apply to storage (reducing the size of model checkpoints) but offer no benefits during inference. While such storage savings are meaningful in large-scale training settings—where frequent snapshotting and checkpoint rollbacks are needed [32]—they have limited impact for everyday LLM end-users. Model downloading is typically a one-time cost, so even if a model checkpoint is compressed by 50%, it only cuts the download time at most by half, presumably over the model’s entire lifecycle of deployment. Furthermore, checkpoints are usually stored on disk, where terabytes of capacity are easily available, making up a much looser constraint compared to GPU HBM (High Bandwidth Memory); one of the main resource constraints during inference.
6 Conclusion
In this work, we present Dynamic-Length Float (DFloat) as a lossless compression data format for LLM weights. To the best of our knowledge, DFloat is the only data format capable of reducing memory footprint while remaining compatible with efficient GPU inference. Specifically, we evaluate several popular LLMs using the 11-bit DFloat format (DF11), alongside custom-developed GPU kernels tailored for this format. Empirical results suggest that DF11-based compression significantly lowers hardware demands for serving LLMs, while introducing reasonable processing overhead in most practical use cases.
Acknowledgements
We thank Oscar Wu for insightful discussions and xMAD.ai for providing computational resources.
Appendix
Appendix A Frequency Distribution of BFloat16 Values
Appendix B Hardware for Experiments
@@
Server 1 NVIDIA Tesla T4 15360MiB Intel Xeon Platinum 8259CL 187GB
Server 2 NVIDIA RTX A5000 24564MiB AMD EPYC 7513 32-Core 504GB
Server 3 NVIDIA A100 40960MiB AMD EPYC 7742 64-Core 1.48TB
Server 4 NVIDIA Quadro RTX 8000 49152MiB AMD EPYC 7742 64-Core 1.48TB
Appendix C GPU Inference Efficiency Comparison: BF16 vs. DF11
We present the GPU inference efficiency of BF16 and DF11 models in Figure 9, for various models, GPUs, and batch sizes. Due to GPU memory constraints, the BF16 models are run with two GPUs while the DF11 models are run with a single GPU.
Figure 9: Comparison of average latency and throughput for token decoding between the original (BF16) models and their losslessly compressed (DF11) counterparts. The DF11 models are run on a single GPU, while the BF16 models require two GPUs due to memory constraints.
Appendix D Impact of Lossy Quantization
Table 5: INT8 quantization error on different tasks. “Math” denotes MATH Hard with 2 shots. “GPQA CoT” is with 2 shots. “$\Delta$” denotes the error gap via INT8 quantization.

### v2

@@ Appendix D Pseudo-code of the GPU kernel for DFloat11 Decompression
47:      $\mathrm{BitOffset}\leftarrow\mathrm{BitOffset}+\mathsf{CodeLengths}[\mathrm{Exponent}]$
48:      $\mathsf{ThreadOutputPos}[t]\leftarrow\mathsf{ThreadOutputPos}[t]+1$
49:     end while
50:    end for
51:      $\triangleright$ Perform coalesced writes to HBM:
52:    $\mathsf{Outputs}[\mathsf{BlockOutputPos}[b]\dots(\mathsf{BlockOutputPos}[b+1]-1)]\leftarrow$
53:      $\mathsf{WriteBuffer}[0\dots(\mathsf{BlockOutputPos}[b+1]-\mathsf{BlockOutputPos}[b]-1)]$
54:  end for
55:end procedure
Appendix E Hardware for Experiments
Table 5 presents the hardware configuration of servers used for experiments.
Appendix F DFloat11 Compression Time
Table 4: Compression time per transformer block for different models.
Model Compression Time per Transformer Block (s)
Llama 3.1 8B Instruct 191
Llama 3.3 70B Instruct 547
Llama 3.1 405B Instruct 2133
Table 4 reports the time required to compress a single transformer block for models of different sizes. Compression is a one-time preprocessing step for each model and is performed using a single CPU thread. Since transformer blocks are independent in terms of weight storage, their compression can be parallelized across multiple CPU threads, making the overall process highly scalable and efficient.
@@ Appendix F DFloat11 Compression Time
Server 1 NVIDIA RTX A5000 24564MiB AMD EPYC 7513 32-Core 504GB
Server 2 NVIDIA A100 40960MiB AMD EPYC 7742 64-Core 1.48TB
Server 3 NVIDIA Quadro RTX 8000 49152MiB AMD EPYC 7742 64-Core 1.48TB
Appendix G GPU Inference Efficiency Comparison: BF16 vs. DF11
We present the GPU inference efficiency of BF16 and DF11 models in Figure 10, for various models and batch sizes on A100 GPUs.
Figure 10: Comparison of average latency and throughput for token decoding between the original (BF16) models and their losslessly compressed (DF11) counterparts. The BF16 and DF11 models are run on the same GPU configurations, with Flash Attention [6] turned on for both methods.
Appendix H Impact of Lossy Quantization
Table 6: INT8 quantization error on different tasks. “Math” denotes MATH Hard with 2 shots. “GPQA CoT” is with 2 shots. $\Delta$” denotes the error gap via INT8 quantization.
@@ Appendix H Impact of Lossy Quantization
Appendix I Efficient Decoding of Huffman Codes Using Compact Lookup Tables
I.1 The Dual Lookup Table Approach
Figure 11: Decoding Huffman codes can be performed either by traversing the Huffman tree or by using two lookup tables: one that maps each $L$-bit binary code to its corresponding symbol, and another that stores the code length for each symbol.
Huffman decoding can be performed by traversing the Huffman tree: starting from the root, each bit of the encoded bitstream determines the branch to follow, and the symbol is fully decoded upon reaching a leaf node. While this bit-by-bit traversal is conceptually simple, it is inefficient in practice. Each branching decision depends on the previous one, leading to frequent memory accesses and conditional jumps. This pattern is especially problematic on GPUs, where it causes branch divergence and limits instruction-level parallelism. A widely adopted alternative is lookup-table-based decoding [44], which flattens the Huffman tree into two compact lookup tables. This enables decoding of each symbol using just two array lookups and a bit shift, significantly improving throughput.
We employ two lookup tables, $\mathsf{LUT}$ and $\mathsf{CodeLengths}$, to achieve efficient, branch-free Huffman decoding. Let $L$ denote the length of the longest codeword in the Huffman codebook. We construct the primary lookup table $\mathsf{LUT}$ as an array of size $2^{L}$, where each entry maps an $L$-bit binary sequence to the first symbol it encodes.
Figure 11 shows an example with $L=4$ and a set of symbols A, B, C, D, E, F. For clarity, we use letters to represent symbols, though in practice these correspond to exponent values in BFloat16 weights. The lookup table $\mathsf{LUT}$ contains $2^{4}=16$ entries, indexed by all possible 4-bit binary sequences. Each entry in $\mathsf{LUT}$ stores the symbol whose Huffman code matches the prefix of that index. If a symbol’s Huffman code is shorter than $L$ bits, it will fill multiple consecutive entries. For example, if symbol A is encoded as the single bit 0, then all binary sequences from 0000 to 0111 begin with 0, so entries 0 through $7$ in $\mathsf{LUT}$ are assigned to A. In contrast, symbols with Huffman codes of length $L$ occupy exactly one entry each. For instance, $\texttt{E}=\texttt{1110}$ and $\texttt{F}=\texttt{1111}$ map to entries $14$ and $15$, respectively. This construction yields a dense prefix table that allows decoding a symbol with a single array lookup using an $L$-bit segment from the encoded bitstream.
To advance the encoded bitstream for decoding the next symbol, we also store the code lengths of all symbols. The second lookup table, $\mathsf{CodeLengths}$, maps each symbol to its Huffman code length. In the example, the lengths are: A:1, B:3, C:3, D:3, E:4, F:4. Together, these two tables allow fast, deterministic decoding by repeating the following steps:
  1. 1.
    Use the next $L$ bits from the encoded bitstream to index $\mathsf{LUT}$ and retrieve the decoded symbol.
  2. 2.
    Look up the code length of the decoded symbol from $\mathsf{CodeLengths}$ to determine how many bits to consume.
  3. 3.
    Advance the encoded bitstream and repeat.
This approach eliminates conditional branches and pointer chasing during decoding, making it highly suitable for parallel computation on GPUs.
I.2 Decomposing $\mathsf{LUT}$ into Hierarchical, Compact Lookup Tables
The primary lookup table $\mathsf{LUT}$ contains $2^{L}$ entries, where $L$ is the maximum code length in the Huffman codebook. While this enables constant-time decoding, the table size grows exponentially with $L$. In practice, $L$ ranges from 24 to 32 for Huffman trees built with BFloat16 exponents. This results in table sizes of $2^{24}$ to $2^{32}$ entries, which far exceeds the capacity of GPU SRAM. To address this, we decompose $\mathsf{LUT}$ into multiple smaller lookup tables that fit within on-chip memory, while still enabling fast decoding.
Hierarchical Table Structure
Figure 12: A Huffman tree can be decomposed into a hierarchy of subtrees, each represented by a compact lookup table (LUT). Each LUT may reference another lower-level LUT in the hierarchy. This hierarchical decoding approach is functionally equivalent to using a single monolithic LUT, but significantly more memory efficient.
Instead of storing a single flat table of size $2^{L}$, we decompose $\mathsf{LUT}$ into a hierarchy of compact lookup tables. Each table corresponds to a subtree of the Huffman tree and is responsible for decoding $b$ bits. Each table processes the next $b$ bits and either (i) directly returns a decoded symbol, or (ii) delegates to a table next in the hierarchy for decoding the next $b$ bits. This hierarchical organization mirrors the structure of the original Huffman tree and significantly reduces total memory usage.
Figure 12 illustrates an example where the Huffman tree is partitioned into three subtrees, each mapped to a separate lookup table responsible for 2 bits. The decoding process using these three LUTs proceeds as follows:
  • $\mathsf{LUT}_{0}$: Uses the first and second bits of the encoded bitstream to determine how to proceed, leading to 3 possible cases:
    • 00, 01 $\rightarrow$ decode the next symbol as A.
    • 10 $\rightarrow$ delegate to $\mathsf{LUT}_{1}$ .
    • 11 $\rightarrow$ delegate to $\mathsf{LUT}_{2}$.
  • $\mathsf{LUT}_{1}$: Uses the third and fourth bits of the encoded bitstream to continue decoding:
    • 00, 01 $\rightarrow$ decode the next symbol as B
    • 10, 11 $\rightarrow$ decode the next symbol as C
  • $\mathsf{LUT}_{2}$: Uses the third and fourth bits of the encoded bitstream to continue decoding:
    • 00, 01 $\rightarrow$ decode the next symbol as D
    • 10 $\rightarrow$ decode the next symbol as E
    • 11 $\rightarrow$ decode the next symbol as F
For decoding Huffman-coded BFloat16 exponents, we decompose the $\mathsf{LUT}$ into multiple compact lookup tables, each responsible for decoding 8 bits (i.e. $b=8$). This allows us to read the next byte from the encoded bitstream and perform an array lookup from a 256-entry array in each step. In practice, the decomposition of $\mathsf{LUT}$ leads to 4 to 8 compact $\mathsf{LUT}$s, each with 256 entries, which comfortably fits within fast SRAM.
Appendix J Text-to-image Results of BF16 and DF11 Diffusion Models
Figure 13: Images generated by Stable Diffusion 3.5 Large in the original BFloat16 precision (top 5) are pixel-wise identical to those produced by the DFloat11-compressed model (bottom 5), using the same prompt and random seed.
Figure 13 presents the comparison of images generated using Stable Diffusion 3.5 Large in BFloat16 and DFloat11 weight format. The images are pixel-wise identical, when using the same prompt and random seed.
Appendix K Limitations
This work focuses exclusively on losslessly compressing BFloat16 weights. We do not consider other formats such as FP32, FP16, or FP8, which may require different compression strategies. While DF11 improves memory efficiency, it introduces a small but non-zero latency overhead due to decompression. This overhead is amortized at larger batch sizes but may impact latency-sensitive applications with small batches. Our evaluation is limited to GPUs. We do not assess performance on other hardware such as CPUs, TPUs, or custom accelerators, which may require platform-specific optimizations.